classes frequency
0-10 4
10-20 8
20-30 7
30-40 10
In this how to draw a frequency polygon?
since as we know that the class previous of the first will -10-0.....please help me to draw this frequency polygon?

Answer by Theo(13342)


you create the frequency polygon by finding the midpoint of each class and then
drawing a graph from it.
the graph would be a straight line between the midpoints on the graph.
first you have to find the midpoints.
midpoint of 0-10 = 5
midpoint of 10-20 = 15
midpoint of 20-30 = 25
midpoint of 30-40 = 35
those are the x-value of your graph.
the y-values of your graph are the frequencies in each class.
so you now have (x,y) coordinate pairs on your graph.
they are:
x       y
5       4
15      8
25      7
35      10

the following reference will show you how to draw the graph.
http://www.math-mate.com/chapter16_10.shtml